OpenStudy (anonymous):

standard form?? y=-5x-7 would be 7=-5x+y right? A cant be negative though...

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

Correct, A cannot be negative :) But let's fix a little mistake first.

OpenStudy (anonymous):

ok, is it the order of the numbers?

zepdrix (zepdrix):

\[\Large\rm y=-5x-7\]To get this into standard form, we'll ADD 5x to each side, yes?

OpenStudy (anonymous):

i though you added the 7

OpenStudy (anonymous):

crap maybe ive been doing all of my problems wrong

zepdrix (zepdrix):

Ok let's try doing that and see what happens.\[\Large\rm y+7=-5x\cancel{-7+7}\]\[\Large\rm y+7=-5x\]

zepdrix (zepdrix):

it's fine to do it this way, it just leaves you with an extra step. See how we now need to move the y to the other side?

OpenStudy (anonymous):

yep!

zepdrix (zepdrix):

Subtracting y from each side gives us:\[\Large\rm 7=-5x-y\]But as you pointed out, your coefficient on the A is negative here. So we need to fix that.

OpenStudy (anonymous):

do we multiply each side by -1

zepdrix (zepdrix):

Just as we can add or subtract a quantity from each side and it stays balanced, we can also multiply, yes good.

OpenStudy (anonymous):

so if we multiplied we would get -7=5x+y or -y?

zepdrix (zepdrix):

We're distributing the -1 to each term on the right. So both the sign on the x and y will change.

zepdrix (zepdrix):

So yes, +y.

zepdrix (zepdrix):

\[\Large\rm y=-5x-7\]Notice that if we has just moved the 5x instead of moving the 7 and y and multiplying by -1, it would have saved us a few steps :) \[\Large\rm y+5x=\cancel{5x-5x}-7\]\[\Large\rm y+5x=-7\]

OpenStudy (anonymous):

ok, that makes it easier! thanks! I just started algebra 2 today and there are a lot of sophomores and juniors in my class @0@

zepdrix (zepdrix):

Are you younger than that? :o

OpenStudy (anonymous):

ya im a freshman

zepdrix (zepdrix):

Haha that's always fun XD Being the young one in the class. You'll look like the smarty pants lol

OpenStudy (anonymous):

hopefully.... I'll probably look really stupid when we actually get to a new topic

zepdrix (zepdrix):

:3

OpenStudy (anonymous):

thanks again!